40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 30 14:31:57 2013, Steve, Update, General, cleaning up MC2areaCP.jpg1Y1areaCleanedUp.jpgETMY_NPROmoved.jpg
    Reply  Mon Sep 2 03:51:14 2013, rana, HowTo, General, How To Coil Cables 
       Reply  Tue Sep 3 10:54:33 2013, Steve, HowTo, General, How To Coil Cables 
Message ID: 9092     Entry time: Fri Aug 30 14:31:57 2013     Reply to this: 9093
Author: Steve 
Type: Update 
Category: General 
Subject: cleaning up 

I cleaned up around MC2 and 1Y1 area this morning.

ETMY NPRO moved from the side to the center of the low shelf. Thorlab catalog " as spacer " removed after lunch.

The Lightwave Controller values: LT 40.4C,  LTEC +0.1V,  T 41.041C,  Pwr 239 mW,  Adj 0,  DC 1.82A,  DPM  0.0V,  Neon OFF,  Ldon OFF,  Display 5,  DT 21.0C,  Dtec +1.0V

 

Attachment 1: MC2areaCP.jpg  277 kB  | Hide | Hide all
MC2areaCP.jpg
Attachment 2: 1Y1areaCleanedUp.jpg  139 kB  | Hide | Hide all
1Y1areaCleanedUp.jpg
Attachment 3: ETMY_NPROmoved.jpg  60 kB  Uploaded Fri Aug 30 17:16:49 2013  | Hide | Hide all
ETMY_NPROmoved.jpg
ELOG V3.1.3-